OpenStudy (anonymous):

solve for x. -3(x+1)-x/2+3/2<0

OpenStudy (jtvatsim):

OK, how far have you gotten toward a solution on your own? Anything in particular that is confusing? :)

OpenStudy (anonymous):

\[-3(x+1)-\frac{ x }{ 2 }+\frac{ 3 }{ 2 }<0\] \[-3x-3-\frac{3+x}{2}<0\] \[-3x-3<\frac{3+x}{2}\] \[-6x-6<3+x\] \[7x>-9\] \[x>-\frac{9}{7}\]